How much grams of water required to pour into 150 cm3 syrup (density of syrup=1,258g/cm3 and procentive concentration 55%) to get the procentive concentration of sucrose 25%?
1
Expert's answer
2020-09-17T07:17:17-0400

Volume of syrup=150ml

Concentration= 55%

How much water added in sryup then it's concentration change 25 %

"V1C1=V2C2"

150×55=V2×25

V2= 150×55/25= 330 ml or Cm3

Water added 330-150= 180 ml

density of water is 1 g/ml

mass of water added = 180 g

Note : ml= Cm3